OpenStudy (anonymous):

Consider a circular drum head or membrane, which may vibrate but is fixed everywhere along its edge. If Y(r, phi, t) describes the motion of the membrane, and the wave equation tells us that (1/v^2) d^2Y/dt^2 = del^2(Y(r, phi)), where v is some constant, then use the method of separation of variables to find the general description of its motion. What is the lowest frequency of vibration?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

So we have the function \(Y(r,\phi,t)\) and also we have: \[\frac{1}{v^2}\frac{d^2Y}{dt^2}=del^2[Y(r,\phi,t)]\]

OpenStudy (anonymous):

Please explain.

OpenStudy (anonymous):

Hmm actually this is much harder than I thought..I am not very well versed in differential equations...

OpenStudy (anonymous):

Still, please help if possible.

OpenStudy (anonymous):

Hmm well I don't know much. But I know that the wave equation can be simplified to: \[\frac{1}{v^2}\frac{d^2Y}{dt^2}=\nabla^2[Y(r,\phi,t)]\] \[\frac{1}{v^2}\frac{d^2Y}{dt^2}=\left(\frac{\partial^2 Y(r,\phi,t)}{\partial r^2},\frac{\partial^2 Y(r,\phi,t)}{\partial \phi^2},\frac{\partial^2 Y(r,\phi,t)}{\partial t^2}\right)\]

OpenStudy (anonymous):

Thanks but can you explain because i have a test tommorow

OpenStudy (anonymous):

I don't know how to solve this problem, but I can go a few more steps in the right direction.

OpenStudy (anonymous):

I would actually be very interested and curious if you would @MrMoose

OpenStudy (anonymous):

The problem commands us to use the separation of variables method. Therefore, we assume: \[Y(r,\phi,t) = X(r,\phi)T(t)\] substituting in: \[\frac{1}{v^2}\frac{d^2T(t)}{dt^2} = \nabla^2X(r,\phi)\]

OpenStudy (anonymous):

Is it clear why this substitution works?

OpenStudy (anonymous):

ie: the derivative with respect to time of a function of position is 0 and vice versa.

OpenStudy (anonymous):

Note that the left hand side is a function of time only and the right hand side is a function of position only.

OpenStudy (anonymous):

Therefore, in order to be equal at all points in time and space, they must be constants.

OpenStudy (anonymous):

... I made a pretty massive typo in the last equation I wrote. It should be: \[\frac{1}{v^2T(t)}\frac{d^2T(t)}{dt^2} = \frac{\nabla^2X(r,\phi)}{X(r,\phi)}\]

OpenStudy (anonymous):

Ask if this isn't clear.

OpenStudy (anonymous):

And sorry for any confusion that caused

OpenStudy (anonymous):

Thanks Obama.

OpenStudy (anonymous):

Now, continuing on: \[\frac{d^2T(t)}{dt^2} = v^2T(t)C\] and \[\nabla^2X(r,\phi) = X(r,\phi)C\]

OpenStudy (anonymous):

I can solve the time equation, but not the position equation.

OpenStudy (anonymous):

Note that if we make the substitution: \[\omega^2 = -v^2C\] we get the harmonic oscillator equation: \[\frac{d^2T(t)}{dt^2} +\omega^2T(t) = 0\]

OpenStudy (anonymous):

@MrMoose can you just explain me what you would do instead of torturing my brain?

OpenStudy (anonymous):

Short answer: I have no idea and I am trying to progress as far as I can and figure it out as I go.

OpenStudy (anonymous):

So anyway, we see that: \[Y(r,\phi,t) = X(r,\phi)(Acos(\omega t)+Bsin(\omega t) )\]

OpenStudy (anonymous):

Or: \[X(r, \phi )Acos(\omega t + \delta)\]

OpenStudy (anonymous):

By the way, what class is this for?

OpenStudy (anonymous):

well im 13 years old but i can do some calc, this is just for fun and i want to prove to some smartass senior in a test that im smarter than him. even without this i would beat him but still.

OpenStudy (anonymous):

I found out the solution to the problem

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Bessel_function

OpenStudy (anonymous):

\[f_0 = \frac{2.4v}{2\pi R}\]

OpenStudy (anonymous):

I think that this is the equation for the fundamental frequency

OpenStudy (anonymous):

Oh now get it thanks bud.

OpenStudy (anonymous):

The description of a radially symmetric wave on a struck drum is: \[Y(r,\phi,t) = J_0(\frac{\omega r}{v})Acos(\omega t)\]

OpenStudy (anonymous):

where J_0 is the 0-order Bessel function